OpenStudy (anonymous):

I'll post this underneath so it makes more sense! (:

OpenStudy (anonymous):

\[(\frac{ 1 }{ 2 } + \frac{ \sqrt{3} }{ 2 }i)^2\]

OpenStudy (anonymous):

I need to put it into standard form

OpenStudy (anonymous):

And somehow I got the same answer, but not squared. :c

OpenStudy (asnaseer):

please show your steps so that I can help spot where you may have made a mistake

OpenStudy (anonymous):

Oh heck, there are a lot of steps. Lool. This might be a while.

OpenStudy (asnaseer):

please use the draw facility if that makes it any easier

OpenStudy (anonymous):

\[= (\frac{ 1 }{ 2 } + \frac{ \sqrt{3} }{ 2 } i)(\frac{ 1 }{ 2 } + \frac{ \sqrt{3} }{ 2 } i) \] \[\frac{ 1 }{ 2 }(\frac{ 1 }{ 2 } + \frac{ \sqrt{3} }{ 2 } i) + \frac{ \sqrt{3} }{ 2 }i(\frac{ 1 }{ 2 } + \frac{ \sqrt{3} }{ 2 } i) \]

OpenStudy (anonymous):

Imma still going.

OpenStudy (asnaseer):

ok... :)

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 } + \frac{ \sqrt{3} }{ 4 } i +\frac{ \sqrt{3} }{ 4 } i +\frac{ 3 }{ 4 } i^2\] \[\frac{ 1 }{ 4 } +\frac{ \sqrt{3} }{ 2 } i + \frac{ 3 }{ 4 } i^2\]

OpenStudy (anonymous):

That last step is where I think I messed up

OpenStudy (asnaseer):

all correct so far :)

OpenStudy (asnaseer):

now do you know what \(i^2=?\)

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 } + \frac{ \sqrt{3} }{ 2 }i -\frac{ 3 }{ 4 }\]

OpenStudy (asnaseer):

perfect! now just one more step left...

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } + \frac{ \sqrt{3} }{ 2 } i\]

OpenStudy (asnaseer):

not quite:\[\frac{1}{4}-\frac{3}{4}\ne\frac{1}{2}\]

OpenStudy (asnaseer):

1 - 3 = ?

OpenStudy (anonymous):

-2

OpenStudy (asnaseer):

do you see where you went wrong now?

OpenStudy (anonymous):

LOL. I always mess up at the end. Thank you so much. (:

OpenStudy (asnaseer):

yw :) - at least you pursued this to the end - which means you learnt something from it! :D

OpenStudy (anonymous):

Veery true! (:

OpenStudy (asnaseer):

:)